generation2006:

solve the system using multiplication for the linear combination method 6x-3y=3 -2x+6y=14 what is the solution system? A.) (2,1) B.) (2,-3) C.) (2,-1) D.) (2,3)

generation2006:

thank you

supie:

\[6x – 3y = 3\] \[–2x + 6y = 14\] \[6x – 3y = 3\] \[–6x + 18y = 42\] \[-3y + 18y = 3 + 42\] \[15y = 45\] \[y = 45/15\] \[y = 3\] \[6x – 3y = 3\] \[6x - 3 * 3 = 3\] \[6x = 3 + 9\] \[6x = 12\] \[x = 12/6\] \[x = 2\] Therefore, the correct answer should be, \[(x, y) = (2, 3)\]